Find the slope of the curve f(x) = (4x - 3)(2x^3 +1) at x = 1
Answer by nerdybill(7384)

Find the slope of the curve f(x) = (4x - 3)(2x^3 +1) at x = 1
.
You will have to apply calculus.
f(x) = (4x - 3)(2x^3 +1)
expand with FOIL:
f(x) = 12x^4 - 6x^3 + 4x - 3
f'(x) = 48x^3 - 18x^2 + 4
.
now, substitute x=1:
f'(1) = 48(1)^3 - 18(1)^2 + 4
f'(1) = 48 - 18 + 4
f'(1) = 34
.
Slope would be 34